OpenStudy (anonymous):

Find the product. x^5 . x^4 . x^3 is it x^12?

mathslover (mathslover):

yes u r right

OpenStudy (theviper):

right :P

mathslover (mathslover):

\[\large{x^5 \times x^4 \times x^3 = x^{5+4+3}=x^{12}}\]

OpenStudy (theviper):

@mathslover is always right ;)